A triangle has side with 5 m, 12 m, and 13 m is it a right triangle
Answer by jim_thompson5910(35256)

Yes it is since

a^2 + b^2 = c^2

5^2 + 12^2 = 13^2

25 + 144 = 169

169 = 169

-------------------------------------------------------

So if you have 3 sides, a,b,c, and they satisfy a^2 + b^2 = c^2 (c is the longest side), then you have a right triangle.